# Danish Fungi 2024 DF20 Danish Fungi 2020

### v2

@@ 5 Conclusion
In this work, we introduced the FungiTastic, a comprehensive and multi-modal dataset and benchmark. The dataset includes a variety of data types, such as photographs, satellite images, climatic data, segmentation masks, and observation metadata. FungiTastic has many interesting features, which make it attractive to the broad machine-learning community. With its data sampling spanning 20 years, precise labels, rich metadata, long-tailed distribution, distribution shifts over time, the visual similarity between the categories, and multimodal nature, it is a unique addition to the existing benchmarks.
In the provided baseline experiments, we demonstrate how challenging the FungiTastic Benchmarks are. Even state-of-the-art architectures and methods yield modest F-scores of 39.8% in closed-set classification and 9.1% in few-shot learning, highlighting the dataset’s difficulty compared to traditional benchmarks such as CUB-200-2011, Stanford Cars, and FGVC Aircraft. The proposed zero-shot baseline for the simplest segmentation task, binary segmentation of fungi fruiting body, achieved an average IoU of 89.36%, which still shows the potential for improvement in fine-grained visual segmentation of fungi.
Limitations mainly lie in the data collection process that influences data distribution. The majority of the data comes from Denmark, and further biases are potentially introduced by "sampling" in locations where certain species may be overrepresented due to their prevalence in frequently sampled areas or collector preferences. Additionally, not all meteorological data are available for every observation, which can affect of multi-modal classification approaches relying on such data.
Future work includes establishing continuous challenges to track progress across varied classification scenarios, expanding the set of weak and strong baselines, and including new test datasets for most benchmarks. Future efforts will also explore additional data, such as traits and textual descriptions of the species, to further improve the dataset’s multi-modal capabilities.
Online Resources
The complexity and size of the dataset require hosting the dataset in multiple spaces. Since the Kaggle dataset is allowing to store datasets just up to around 140GB we provide images for the vanilla FungiTastic dataset in smaller resolution. The full-size images, as well as the Croissant metadata, are available for download through dataset documentation.
Apart from the dataset, we also provide a wide variety of assets. Notably, we provide:
  • FGVC, a Pytorch-based framework designed for training deep learning classification models, hyperparameter fine-tuning, Weights&Biases logging, ONNX model export and HuggingFace deployment.
  • Sample notebooks that allows loading pre-trained models, prediction and making submission on Kaggle benchmarks.
  • All baseline pre-trained models available through the HuggingFace model hub allowing a straightforward integration and usability.
  • A comprehensive documentation with tables data samples and training strategies.
@@ Acknowledgement
We extend our sincere gratitude to the mycologists from the Danish Mycological Society, particularly Jacob Heilmann-Clausen, Thomas Læssøe, Thomas Stjernegaard Jeppesen, Tobias Guldberg Frøslev, Ulrik Søchting, and Jens Henrik Petersen, for their invaluable contributions and expertise. We also thank the dedicated citizen scientists whose data and efforts have been instrumental to this project. Your support and collaboration have greatly enriched our work and made this research possible. Thank you for your commitment to advancing ecological understanding and conservation.
@@ Appendix A Evaluation Metrics
A.1 Closed and open set classification
The main evaluation metric is $F$, the macro-averaged F1-score. For closed-set classification, the evaluation is standard, and for open-set, it is defined as
$F=\frac{1}{{C}}\sum_{c=1}^{C}{F}_{c},\quad{F}_{c}=\frac{2{P}_{c}\cdot{R}_{c}}{{P}_{c}+{R}_{c}},$ (1)
where ${P}_{c}$ and ${R}_{c}$ are the recall and precision of class $c$ and ${C}$ is the total number of classes, including the unknown class $u$.
@@ A.1 Closed and open set classification
where ${K}$ = $\{1\dots C\}\setminus\{u\}$ is the set of known classes. The ${F}_{K}$ also corresponds to the main evaluation metric for standard closed-set classification. Additional metrics reported are top-1 and top-3 accuracy, defined as
${Acc@k}=\frac{1}{{N}}\sum_{i=1}^{N}\mathbf{1}\left(y_{i}\in q_{k}(x_{i})\right),$ (3)
A.2 Classification beyond 0-1 loss function
For the classification beyond 0-1 cost, we follow the definition we set for the annual FungiCLEF competition [37, 38]. A metric of the following general form should be minimized.
@@ A.2 Classification beyond 0-1 loss function
$\text{W}^{p/e}(y,q_{1}(x))=\begin{cases}0&\text{if }d(y)=d(q_{1}(x))\\ c_{p}&\text{if }d(y)=1\text{ and }d(q_{1}(x))=0,\\ c_{e}&\text{otherwise}\end{cases}$ (5)
where $d(y),y\in\text{C}$ is a binary function that indicates dangerous (poisonous) species ($d(y)=1$), $c_{p}=100$ and $c_{e}=1$. For the known/unknown species scenario, we define the cost matrix as
$\text{W}^{k/u}(y,q_{1}(x))=\begin{cases}0&\text{if }y=q_{1}(x)\\ c_{u}&\text{if }y=u\text{ and }q(x)\neq u,\\ c_{k}&\text{otherwise}\end{cases}$ (6)
where $c_{u}=10$ and $c_{k}=1$.
@@ Appendix B Supporting Figures and Tables
Table 6: FungiTastic dataset splits – statistical overview. We provide the number of observations, images, and classes for each benchmark and the corresponding dataset. "Unknown classes" are those with no available data in training. DNA stands for DNA-sequenced data.
Dataset Subset Observations Images Classes Unknown classes
@@
Microscopic
FungiTastic – Closed Set Train. 246,884 433,701 2,829
Val. 45,616 89,659 2,306
Test. 48,379 91,832 2,336
DNA 2,041 5,117 725
FungiTastic–M – Closed Set Train. 25,786 46,842 215
Val. 4,687 9,412 193
Test. 5,531 10,738 196
DNA 211 645 93
FungiTastic–FS – Closed Set Train. 4,293 7,819 2,427
Val. 1,099 2,285 570
Test. 998 1,909 566
FungiTastic – Open Set Train. 246,884 433,701 2,829
Val. 47,453 96,756 3,360 1,054
Test. 50,085 97,551 3,349 1,013
Figure 9: Additional samples of fruiting body part segmentation.

### v3

@@ 6 Conclusion
In this work, we introduced the FungiTastic, a comprehensive and multi-modal dataset and benchmark. The dataset includes a variety of data types, such as photographs, satellite images, climatic data, segmentation masks, and observation metadata. FungiTastic has many interesting features, which make it attractive to the broad ML community. With its data sampling spanning 20 years, precise labels, rich metadata, long-tailed distribution, distribution shifts over time, the visual similarity between the categories, and multimodal nature, it is a unique addition to the existing benchmarks.
In the provided baseline experiments, we demonstrate how challenging the FungiTastic Benchmarks are. Even state-of-the-art architectures and methods yield modest F-scores of 39.8% in closed-set classification and 9.1% in few-shot learning, highlighting the dataset’s challenging nature compared to traditional benchmarks such as CUB-200-2011, Stanford Cars, and FGVC Aircraft. The proposed zero-shot baseline for the simplest segmentation task, binary segmentation of fungi fruiting body, achieved an average IoU of 89.36%, which still shows the potential for improvement in fine-grained visual segmentation of fungi. The open-set baselines show that discovering novel classes remains a difficult task, demanding new techniques tailored to fine-grained recognition. Additionally, results with non-domain-specific vision-language models reveal a surprisingly strong performance of such models. The fusion experiments of VLMs with supervised models confirm the challenge of accurately classifying rare species in highly imbalanced datasets.
Limitations lie in the data collection process, which affects the overall distribution. Most of the data comes from Denmark, and bias is further introduced through "random" sampling. Therefore, some species are more common in frequently sampled areas or are favored by collectors. Some recent observations also miss metadata, which can reduce the effectiveness of classification methods that rely on it.
Future work includes setting up and running future challenges [31], expanding baseline models, adding new test sets, and exploring extra data like traits and species descriptions to enhance multi-modal performance.
@@ Acknowledgement
This research was supported by the Technology Agency of the Czech Republic, project No. SS73020004. We extend our sincere gratitude to the mycologists from the Danish Mycological Society, particularly Jacob Heilmann-Clausen, Thomas Læssøe, Thomas Stjernegaard Jeppesen, Tobias Guldberg Frøslev, Ulrik Søchting, and Jens Henrik Petersen, for their contributions and expertise. We also thank the dedicated citizen scientists whose data and efforts have been instrumental to this project. Your support and collaboration have greatly enriched our work and made this research possible. Thank you for your commitment to advancing ecological understanding and conservation.
@@ Appendix A Evaluation Metrics
The diversity and unique features of the FungiTastic dataset allow for the evaluation of various fundamental computer vision and machine learning problems. We present several distinct benchmarks, each with its own evaluation protocol. This section provides a detailed description of all evaluation metrics for each benchmark.
A.1 Closed set classification
For closed-set classification, the main evaluation metric is $\text{F}_{1}^{m}$, i.e., the macro-averaged F1-score, defined as
$\text{F}_{1}^{m}=\frac{1}{{C}}\sum_{c=1}^{C}{F}_{c},\quad{F}_{c}=\frac{2{P}_{c}\cdot{R}_{c}}{{P}_{c}+{R}_{c}},$ (1)
where ${P}_{c}$ and ${R}_{c}$ are the recall and precision of class $c$ and ${C}$ is the total number of classes. Additional metrics of interest are Recall@k, defined as
$\text{Recall}@k=\frac{1}{{N}}\sum_{i=1}^{N}\mathbf{1}\left(y_{i}\in q_{k}(x_{i})\right),$ (2)
A.2 Few-shot classification
The few-shot classification challenge does not have any unknown classes and can be considered as closed-set classification. Unlike other FungiTastic subsets, the few-shot subset does not suffer from high class imbalance and we choose the Top1 accuracy as the main metric. F1 score and Top3 accuracy are also reported. All metrics are as defined in closed-set classification.
A.3 Open-set classification
The primary metric used for evaluation is the Receiver Operating Characteristic Area Under the Curve (ROCAUC), which measures the ability of the model to distinguish between classes across various threshold values. ROCAUC is defined as the area under the ROC curve, which plots the True Positive Rate (TPR) against the False Positive Rate (FPR) at different classification thresholds where
TPR=True Positives (TP)True Positives (TP)+False Negatives ($\small\text{TPR}=\frac{\text{True Positives (TP)}}{\text{True Positives (TP)}+\text{False Negatives (FN)}},$ (3)
FPR=False Positives (FP)False Positives (FP)+True Negatives ($\small\text{FPR}=\frac{\text{False Positives (FP)}}{\text{False Positives (FP)}+\text{True Negatives (TN)}}.$ (4)
In addition to ROCAUC, the True Negative Rate (TNR) at 95% TPR (TNR95) is also reported. The TNR, also known as specificity, is defined as:
TNR=True Negatives (TN)True Negatives (TN)+False Positives ($\small\text{TNR}=\frac{\text{True Negatives (TN)}}{\text{True Negatives (TN)}+\text{False Positives (FP)}}.$ (5)
The TNR95 metric indicates the specificity achieved when the True Positive Rate (TPR) is fixed at 95%, reflecting the model’s ability to minimize false positives while maintaining a high sensitivity.
@@ A.3 Open-set classification
where ${K}$ = $\{1\dots C\}\setminus\{u\}$ is the set of known classes.
A.4 Classification beyond 0-1 loss function
For the classification beyond 0-1 cost, we follow the definition we set for the annual FungiCLEF competition. A metric of the following general form should be minimized.
@@ A.4 Classification beyond 0-1 loss function
$\small\text{W}^{p/e}(y,q_{1}(x))=\begin{cases}0&\text{if }d(y)=d(q_{1}(x))\\ c_{p}&\text{if }d(y)=1\text{ and }d(q_{1}(x))=0,\\ c_{e}&\text{otherwise}\end{cases}$ (8)
where $d(y),y\in\text{C}$ is a binary function that indicates dangerous (poisonous) species ($d(y)=1$), $c_{p}=100$ and $c_{e}=1$.
A.5 Segmentation
Provided segmentation masks allow the evaluation of many different segmentation scenarios; here, we highlight two.
Binary segmentation, where the positive class is the foreground (mushroom) and the negative class is the background (the complement). The metric is the intersection-over-union (IoU) averaged over all images ($\text{IoU}_{\text{B}}$), giving each image the same weight
$\text{IoU}_{\text{B}}=\frac{1}{I}\sum_{i=1}^{I}\frac{|P_{i}\cap G_{i}|}{|P_{i}\cup G_{i}|},$ (9)
where $I$ is the total number of images, $P_{i}$ is the predicted set of foreground pixels for image $i$, $G_{i}$ is the ground truth set of foreground pixels for image $i$, $|P_{i}\cap G_{i}|$ is the intersection (true positives) for image $i$ and $|P_{i}\cup G_{i}|$ is the union (true positives + false positives + false negatives) for image $i$.
For semantic segmentation, we adopt the standard mean intersection-over-union (mIoU) metric, where per-class IoUs are averaged, giving each class the same weight
$\text{mIoU}=\frac{1}{C}\sum_{c=1}^{C}\frac{|P_{c}\cap G_{c}|}{|P_{c}\cup G_{c}|},$ (10)
where $C$ is the total number of classes, $P_{c}$ is the predicted set of pixels for class $c$, $G_{c}$ is the ground truth set of pixels for class $c$, $|P_{c}\cap G_{c}|$ is the intersection (TPs) and $|P_{c}\cup G_{c}|$ is the union (TPs + FPs + FNs).
@@ Appendix B Supporting Figures and Tables
B.1 Closed-set experiment with higher input size
Following the results provided in the paper, we further experimented with how input size affects classification performance. Switching from 224$\times$224 to 384$\times$384 increased the performance by around five percentage points in all measured metrics and for almost all the architectures. Still, the best-performing model, i.e., BEiT-Base/p16 achieves "just" 75% accuracy and less then 50% in terms of $\text{F}_{1}^{m}$.
Table 8: Closed-set fine-grained classification FungiTastic and FungiTastic–M. A set of selected state-of-the-art CNN- and Transformer-based architectures evaluated on the test sets. All reported metrics show the challenging nature of the dataset.
FungiTastic–M – 3842 FungiTastic – 3842
Architectures Top1 Top3   $\text{F}_{1}^{m}$ Top1 Top3   $\text{F}_{1}^{m}$
ResNet-50 66.3 82.9 39.8 66.9 80.9 36.3
ResNeXt-50 67.0 84.0 39.9 68.1 81.9 37.5
EfficientNet-B3 67.4 82.8 40.5 68.2 81.9 37.2
EfficientNet-v2-B3 70.3 85.8 43.9 72.0 84.7 41.0
ConvNeXt-Base 70.2 85.7 43.9 70.7 83.8 39.6
ViT-Base/p16 73.9 87.8 46.3 74.9 86.3 43.9
Swin-Base/p4w12 72.9 87.0 47.1 74.3 86.3 43.3
BEiT-Base/p16 74.8 88.3 48.5 75.3 86.7 44.5
B.2 FungiTastic – Dataset statistics
The FungiTastic dataset offers a rich and diverse collection of observations and metadata. To provide a clearer understanding of its scope, Table 9 presents a statistical overview of its subsets, including the number of observations, associated images, species categories, and metadata availability. Each subset caters to specific benchmarking needs, ensuring comprehensive evaluation scenarios.
Table 9: FungiTastic dataset splits – statistical overview. The number of observations, images, and classes for each benchmark and the corresponding dataset. "Unknown classes" are those with no available data in training. DNA stands for DNA-sequenced data.
Dataset Subset Observ. Images Classes Unkn.
@@ B.2 FungiTastic – Dataset statistics
Captions
FungiTastic Train. 246,884 433,701 2,829
Closed Set Val. 45,613 89,659 2,306
Test 48,378 91,832 2,336
[…]
FungiTastic–M Train. 25,786 46,842 215
Closed Set Val. 4,687 9,412 193
Test 5,531 10,738 196
[…]
FungiTastic–FS Train. 4,293 7,819 2,427
Closed Set Val. 1,099 2,285 570
Test 999 1,911 567
FungiTastic Train. 246,884 433,702 2,829
Open Set Val. 47,450 96,756 3,360 1,053
Test 50,084 97,551 3,349 1,000
Total unique values: 349,307 632,313 6,034 1,678
B.3 Additional figures
To further highlight the unique features of the FungiTastic dataset, we provide additional figures. These include: (i) a time series sample of temperature data illustrating climatic variability over a 3-year period (Figure 8), (ii) examples of detailed text descriptions generated for individual images to aid in species identification (Figure 9), and (iii) visual samples of ground truth segmentations that highlight different fruiting body parts of fungi (Figure 10).
Figure 8: Sample of available temperature data. First 3 years (out of 20 available years) of average, maximum, and minimum monthly temperatures for a randomly selected location.
Refer to caption The image shows two mushrooms growing on a dark, decaying log. The mushroom on the left is larger and more prominent. It has a light brown cap with a slightly darker brown rim. The cap’s surface appears smooth and slightly shiny. Its stem is thick and light brown, with a hint of green at the base. The smaller mushroom on the right has a similar light brown cap, but its rim is more pronounced and has a white, almost translucent appearance. This gives it a delicate, lacy look. The stem of this mushroom is thinner and lighter in color compared to its larger companion. Both mushrooms have a gilled structure visible underneath their caps. The gills are light brown ……
Refer to caption The image shows a cluster of mushrooms growing on a decaying log. The fungi have distinctive features that are clearly visible: The caps of the mushrooms are dark brown in color, with a slightly lighter shade around the edges. They have a dome-like shape, resembling small umbrellas or bell-shaped structures. The gills of the mushrooms are visible underneath the caps. They appear as thin, white, vertical lines radiating from the stem to the edge of the cap. These gills are a key identifying feature of many mushroom species. The stems of the mushrooms are light brown in color and appear to be relatively thick and sturdy. They support the caps and extend into ……
Refer to caption The image shows four mushrooms arranged in a square formation on a grassy surface. These fungi have distinctive visual features that are clearly visible: The caps of the mushrooms are dark brown in color, with a smooth, slightly shiny appearance. They have a rounded, dome-like shape that’s typical of many mushroom species. The stems of the mushrooms are white, providing a stark contrast to the dark caps. They appear to be relatively thick and sturdy, supporting the caps firmly. The texture of the caps seems to be relatively smooth, without any visible ridges or bumps. The stems appear to have a slightly fuzzy or velvety texture. In terms of size, the mushrooms are of ……
Refer to caption The image shows a single mushroom lying on its side in a grassy area. The mushroom has a distinctive shape with a white cap and a long, cylindrical stem. The cap is rounded and slightly curved, while the stem is straight and elongated. The mushroom’s surface appears smooth and uniform in color, with a clean white appearance. The cap and stem are of similar width, creating a balanced look. The stem extends from the base of the cap and continues downward, with the tip pointing towards the bottom right corner of the image. The mushroom is covered in dirt, particularly on its stem and parts of the cap. This gives it a freshly harvested appearance and ……
Refer to caption The image shows a large, flat mushroom as the central focus. Its cap is predominantly gray with a slightly mottled appearance, featuring some darker patches. The surface of the cap appears smooth and slightly shiny, reflecting light in certain areas. The mushroom’s cap is circular in shape, though it’s not perfectly round. It has a slightly irregular edge, giving it a natural, organic look. The cap’s size is quite substantial, dominating the frame of the image. On the top of the mushroom cap, there are several small, white, circular structures. These are likely the gills or spore-bearing areas of the mushroom. They stand out in contrast against the gray cap, ……
Refer to caption The image shows two mushrooms growing on a dark, decaying log. The mushroom on the left is larger and more prominent. It has a light brown cap with a slightly darker brown rim. The cap’s surface appears smooth and slightly shiny. Its stem is thick and light brown, with a hint of green at the base. The smaller mushroom on the right has a similar light brown cap, but its rim is more pronounced and has a white, almost translucent appearance. This gives it a delicate, lacy look. The stem of this mushroom is thinner and lighter in color compared to its larger companion. Both mushrooms have a gilled structure visible underneath their caps. The gills are light brown ……
Figure 9: Additional image caption samples. For each photograph, we provide a Malmo-7B image caption-like text description.
Figure 10: Additional samples of ground truth fruiting body part segmentation.